OpenStudy (anonymous):

The equation of a line is y = 2x + 3. What is the equation of the line that is parallel to the first line and passes through (2, –1)? A. 4x – 2y = –6 B. y = 2x – 5 C. y = 3x + 4 D. 2x + y = –1

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@Jamierox4ev3r

OpenStudy (samgrace):

B. Using point (2,-1) with the same gradient as the first equation (as they're parallel). Then plug into formula y - y1 = m(x - x1)

OpenStudy (samgrace):

y - (-1) = 2 (x -2) y + 1 = 2x - 4 ==> y = 2x - 5

OpenStudy (anonymous):

can you help me with another one? @samgrace

OpenStudy (samgrace):

Yes

OpenStudy (anonymous):

The graph of which of the following will be parallel to the graph of 4x – 3y = –12?